OpenStudy (anonymous):

Find two unit vectors orthogonal to both (8, 9, 1) and (−1, 1, 0)

OpenStudy (amistre64):



OpenStudy (anonymous):

when two vectors are othogonal, their dot product must be "0" let the vector be \(\vec{u}=(a,b,c)\)

OpenStudy (anonymous):

then \[ {\rm 1:}\quad (8,9,1)\cdot(a,b,c)=0\\ 8a+9b+c=0 \] similarly, for the second one. then it is given that \(\vec{u}\) is a unit vector. so, we get "3" equations in a, b, and c

OpenStudy (amistre64):

\[\pm\frac{\vec u \times \vec v}{|\vec u \times \vec v|}\]